OpenStudy (anonymous):

in an isolated town of 5000 inhabitants,160 people have a disease at the beginning of the week and 1200 have it at the end of the week. How long does it take for 80% of the population to become infected?

OpenStudy (anonymous):

80% of 5000 = 4000 160=beginning of the week(1day) 1200=end of the week (7days) first 1200X4=4800 4week(28days) 160 + 4800 = 4960 too small 160 X 2 =320 (2days) 320 + 4800 = 5120 4 weeks and 2 days or (30days)

OpenStudy (anonymous):

uh i gotta use differential equations to do it

OpenStudy (anonymous):

what kind?

OpenStudy (anonymous):

Ae^kt that one

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

sorry man